(Section 11.4) Find v x u for: u= i +6 j = (-2,1,1)
Answer #1

n = i +63 V-K-2,1,1)=-2 / +)tk i k ūru -2 1 1 6 i(0-6) - ) 10-12 + (-12-1) =-6i +/-13k ū x M <-G,1,-13) Git)-13k 2